OpenStudy (anonymous):

You receive an email that asks you to forward the email to 5 people. Each of those 5 people then forwards the email to 5 people, and so on. What are the next five terms: 1 + ....

OpenStudy (jim766):

1 5 each of these 5 email to 5 people 25 each of these 25 email to 5 people 125 each of these 125 email to 5 people , can you do it from here?

OpenStudy (anonymous):

i think it is 6, 16, 31, 51, 76 is that right?

OpenStudy (anonymous):

no nevermind im not sure

OpenStudy (jim766):

do you follow what I have written?

OpenStudy (anonymous):

im kinda confused

OpenStudy (jim766):

1 person sends and email to five people 1*5 = 5 5 if each of these 5 people send an email to 5 people 5 *5 = 25 25 if each of these 5 people send an email to 5 people 25 *5 = 125 125 what goes next?

OpenStudy (anonymous):

625?

OpenStudy (jim766):

right! then what? you just need 1 more

OpenStudy (anonymous):

3125 right?

OpenStudy (jim766):

right!

OpenStudy (anonymous):

okay! thank you so much!